# LiveSearchBench: An Automatically Constructed Benchmark for Retrieval and Reasoning over Dynamic Knowledge

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable progress across diverse natural language processing tasks, with solid performance on prominent search question answering (QA) benchmarks such as Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and HotpotQA (Yang et al., 2018). Recent reinforcement learning (RL) methods have further improved headline performance, strengthening the perception that LLMs possess sophisticated reasoning and knowledge-intensive inference capabilities (Jin et al., 2025b; Fan et al., 2025). However, a fundamental limitation persists: most search-oriented benchmarks are static and outdated. Many were collected years ago, raising the risk that answers are encoded in models’ parametric memory due to pre-training contamination rather than discovered via retrieval (Wu et al., 2025).
@@
World knowledge is inherently dynamic—news breaks, software versions change, policies evolve, and social events unfold—yet prevailing benchmarks lack mechanisms to incorporate real-time updates. Because of this static nature, evaluating retrieval on these datasets is unreliable: models can often answer questions without invoking any search, relying solely on internal memory. As emphasized by the notion of a Knowledge Boundary (Wang et al., 2025; Chen et al., 2025b), there is a critical distinction between what a model remembers and what it must acquire externally. Our preliminary experiments corroborate this concern: several models achieve strong scores even when retrieval is disabled, suggesting that memorized knowledge dominates and obscures true capacity for acquiring and reasoning over up-to-date external information.
@@
To contextualize the evolution of QA evaluation and retrieval-centric resources, Figure 1 highlights key datasets and model milestones. As the timeline shows, many widely used benchmarks predate recent advances in search-integrated inference, and community efforts have largely prioritized model development over evaluation under dynamic conditions. Motivated by these gaps—and inspired by the contamination-aware practices of LiveCodeBench (Jain et al., 2024)—we introduce LiveSearchBench, a continually updated benchmark built via a scalable pipeline that synthesizes questions from real-world editing streams. The benchmark remains fresh and temporally grounded, with validation that enforces both factual correctness and temporal consistency. By design, success hinges on up-to-date retrieval rather than parametric recall, moving beyond memory-based performance on static snapshots. Our evaluation yields three key insights. First, by systematically testing LLMs and retrieval-augmented generation (RAG) systems on LiveSearchBench, we expose marked differences in their ability to handle dynamic knowledge. Second, we observe a persistent gap between memorization-driven responses and genuine retrieval-based inference. Third, these findings underscore the need for benchmarks that reflect realistic, time-sensitive conditions. This paper makes the following contributions:
@@ 2 Related Work
LLMs leverage external knowledge along three complementary axes (Zhang et al., 2025). (i) Retrieval-augmented generation (RAG). RAG has become a prevailing strategy for grounding outputs in external evidence, and recent surveys consolidate design choices and best practices Gao et al. (2024); Fan et al. (2024). (ii) Workflow-style search agents. Agentic systems explicitly plan queries, browse sources, verify snippets, and synthesize answers; recent efforts integrate these steps into inference-time reasoning traces, exemplified by Search-o1 (Li et al., 2025c). (iii) Reinforcement learning for search and reasoning. RL improves query formulation and the coordination between search and reasoning by directly optimizing end-to-end behavior on challenging objectives (Jin et al., 2025b; Fan et al., 2025; Sun et al., 2025; Song et al., 2025; Chen et al., 2025a). These lines differ in where the search policy resides and how evidence is injected, yielding complementary avenues for strengthening LLMs’ use of external knowledge.
@@
Single-hop search QA is widely evaluated using Natural Questions, TriviaQA, and SimpleQA (Kwiatkowski et al., 2019; Joshi et al., 2017; Wei et al., 2024). Multi-hop reasoning is assessed by HotpotQA, 2WikiMultihopQA, MuSiQue, Bamboogle, and BrowseComp (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2022; Press et al., 2022; Wei et al., 2025). While SimpleQA and BrowseComp incorporate careful curation and adversarial design, all these resources remain static snapshots. This limits scalability, risks overlap with pre-training corpora, and provides weak coverage of time-sensitive knowledge. In parallel, recent work constructs synthetic, web-grounded data for training (Tao et al., 2025; Li et al., 2025a), —aimed at scaling instruction-tuning or corpus quality rather than evaluation; such pipelines generally do not enforce temporal recency, uniqueness guarantees, or machine-verifiable provenance.
@@
In the experiment, parametric-only systems can match or even exceed retrieval-augmented pipelines on static datasets—without accessing external evidence. Retrieval is not uniformly beneficial, especially on multi-hop datasets, where it can introduce noise and compound errors. Additionally, different model families exhibit consistent offsets as $k$ increases. Our experiments show that pass@k accuracy is relatively high even without retrieval, suggesting that with reinforcement learning (RL) techniques, the pass@$k$ score could potentially converge towards pass@1, further closing the gap and possibly surpassing retrieval@1 (Fan et al., 2025; Guo et al., 2025). These patterns validate our hypothesis: static benchmarks overestimate an LLM’s ability to handle dynamic, time-sensitive knowledge. They reward distributional familiarity and sampling strategy, rather than the need for up-to-date evidence, underscoring the necessity for benchmarks whose questions explicitly depend on current, verifiable sources.
@@ 5.1 Experimental Setup
We group baselines into three categories. Vanilla Prompt Methods include Direct Prompt and Chain-of-Thought (CoT) prompting to elicit structured reasoning without external evidence. RAG-based Methods comprise standard retrieval-augmented generation and Search-o1 (Li et al., 2025b). RL-based Methods include Search-R1 (Jin et al., 2025a), and SSRL (Fan et al., 2025). To ensure a fair comparison in online settings, the number of retrieved passages is capped at $3$ across all RAG-style approaches. For vanilla prompt methods, we employ instruction-tuned variants because they exhibit stronger prompt-following behavior. Full implementation details, hyperparameters and some other baselines are provided in Appendix D and code repo.
@@ Appendix B Other Related Work
Complementary to our focus on dynamic, retrieval-dependent QA with SPARQL-verifiable answers, recent work advances reliable long-horizon reasoning along orthogonal axes. Multi-agent methods (Reso, Swarm) improve internal coordination and exploration but do not test temporal grounding on evolving facts (Zhou et al., 2025b; Zhu et al., 2025). RL post-training reveals scaling laws for mathematical reasoning, yet its gains must transfer under recency-sensitive evaluation (Tan et al., 2025). In multimodal and embodied settings, Visual-O1 and BMMR target ambiguity and cross-discipline reasoning, while BEAR and Code-as-Monitor emphasize verifiable skills and failure detection (Ni et al., 2024; Xi et al., 2025; Qi et al., 2025; Zhou et al., 2024; Kang et al., 2025). RoboRefer strengthens spatial referring via structured reasoning (Zhou et al., 2025a). VeriGUI contributes stepwise verification for GUI agents (Liu et al., 2025). LiveSearchBench uniquely couples compositional constraints with fresh, machine-checkable knowledge to assess reasoning beyond parametric memory.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable progress across diverse natural language processing tasks, with solid performance on prominent search question answering (QA) benchmarks such as Natural Questions (kwiatkowski-etal-2019-natural), TriviaQA (joshi-etal-2017-triviaqa), and HotpotQA (yang-etal-2018-hotpotqa). Recent reinforcement learning (RL) methods have further improved headline performance, strengthening the perception that LLMs possess sophisticated reasoning and knowledge-intensive inference capabilities (jin2025searchr; fan2025ssrl). However, a fundamental limitation persists: most search-oriented benchmarks are static and outdated. Many were collected years ago, raising the risk that answers are encoded in models’ parametric memory due to pre-training contamination rather than discovered via retrieval (wu2025reasoningmemorizationunreliableresults).
@@
World knowledge is inherently dynamic—news breaks, software versions change, policies evolve, and social events unfold—yet prevailing benchmarks lack mechanisms to incorporate real-time updates. Because of this static nature, evaluating retrieval on these datasets is unreliable: models can often answer questions without invoking any search, relying solely on internal memory. As emphasized by the notion of a Knowledge Boundary (wang2025toward; chen2025decisionflowadvancinglargelanguage), there is a critical distinction between what a model remembers and what it must acquire externally. Our preliminary experiments corroborate this concern: several models achieve strong scores even when retrieval is disabled, suggesting that memorized knowledge dominates and obscures true capacity for acquiring and reasoning over up-to-date external information.
@@
To contextualize the evolution of QA evaluation and retrieval-centric resources, Figure 1 highlights key datasets and model milestones. As the timeline shows, many widely used benchmarks predate recent advances in search-integrated inference, and community efforts have largely prioritized model development over evaluation under dynamic conditions. Motivated by these gaps—and inspired by the contamination-aware practices of LiveCodeBench (jain2024livecodebenchholisticcontaminationfree)—we introduce LiveSearchBench, a continually updated benchmark built via a scalable pipeline that synthesizes questions from real-world editing streams. The benchmark remains fresh and temporally grounded, with validation that enforces both factual correctness and temporal consistency. By design, success hinges on up-to-date retrieval rather than parametric recall, moving beyond memory-based performance on static snapshots. Our evaluation yields three key insights. First, by systematically testing LLMs and retrieval-augmented generation (RAG) systems on LiveSearchBench, we expose marked differences in their ability to handle dynamic knowledge. Second, we observe a persistent gap between memorization-driven responses and genuine retrieval-based inference. Third, these findings underscore the need for benchmarks that reflect realistic, time-sensitive conditions. This paper makes the following contributions:
@@ 2 Related Work
LLMs leverage external knowledge along three complementary axes (zhang2025landscapeagenticreinforcementlearning). (i) Retrieval-augmented generation (RAG). RAG has become a prevailing strategy for grounding outputs in external evidence, and recent surveys consolidate design choices and best practices gao2024retrievalaugmentedgenerationlargelanguage; fan2024surveyragmeetingllms. (ii) Workflow-style search agents. Agentic systems zhou2025reso; qinmp5; qin2025robofactory explicitly plan queries, browse sources, verify snippets, and synthesize answers; recent efforts integrate these steps into inference-time reasoning traces, exemplified by Search-o1 (search-o1). (iii) Reinforcement learning for search and reasoning. RL improves query formulation and the coordination between search and reasoning by directly optimizing end-to-end behavior on challenging objectives (jin2025searchr; fan2025ssrl; sun2025zerosearchincentivizesearchcapability; song2025r1searcherincentivizingsearchcapability; chen2025researchlearningreasonsearch; huang2025reinforcedinternalexternalknowledgesynergistic). These lines differ in where the search policy resides and how evidence is injected, yielding complementary avenues for strengthening LLMs’ use of external knowledge.
@@
Single-hop search QA is widely evaluated using Natural Questions, TriviaQA, and SimpleQA (kwiatkowski-etal-2019-natural; joshi-etal-2017-triviaqa; wei2024measuring). Multi-hop reasoning is assessed by HotpotQA, 2WikiMultihopQA, MuSiQue, Bamboogle, and BrowseComp (yang-etal-2018-hotpotqa; ho-etal-2020-constructing; trivedi2022musiquemultihopquestionssinglehop; press2022measuring; wei2025browsecomp). While SimpleQA and BrowseComp incorporate careful curation and adversarial design, all these resources remain static snapshots. This limits scalability, risks overlap with pre-training corpora, and provides weak coverage of time-sensitive knowledge. In parallel, recent work constructs synthetic, web-grounded data for training (tao2025webshaper; li2025websailornavigatingsuperhumanreasoning), —aimed at scaling instruction-tuning or corpus quality rather than evaluation; such pipelines generally do not enforce temporal recency, uniqueness guarantees, or machine-verifiable provenance.
@@
In the experiment, parametric-only systems can match or even exceed retrieval-augmented pipelines on static datasets—without accessing external evidence. Retrieval is not uniformly beneficial, especially on multi-hop datasets, where it can introduce noise and compound errors. Additionally, different model families exhibit consistent offsets as $k$ increases. Our experiments show that pass@k accuracy is relatively high even without retrieval, suggesting that with reinforcement learning (RL) techniques, the pass@$k$ score could potentially converge towards pass@1, further closing the gap and possibly surpassing retrieval@1 (fan2025ssrl; guo2025deepseek). These patterns validate our hypothesis: static benchmarks overestimate an LLM’s ability to handle dynamic, time-sensitive knowledge. They reward distributional familiarity and sampling strategy, rather than the need for up-to-date evidence, underscoring the necessity for benchmarks whose questions explicitly depend on current, verifiable sources.
@@ 5.1 Experimental Setup
We group baselines into three categories. Vanilla Prompt Methods include Direct Prompt and Chain-of-Thought (CoT) prompting to elicit structured reasoning without external evidence. RAG-based Methods comprise standard retrieval-augmented generation and Search-o1 (li2025searcho1agenticsearchenhancedlarge). RL-based Methods include Search-R1 (jin2025searchr1trainingllmsreason), and SSRL (fan2025ssrl). To ensure a fair comparison in online settings, the number of retrieved passages is capped at $3$ across all RAG-style approaches. For vanilla prompt methods, we employ instruction-tuned variants because they exhibit stronger prompt-following behavior. Full implementation details, hyperparameters and some other baselines are provided in Appendix D and code repo.
@@ Appendix B Other Related Work
Complementary to our focus on dynamic, retrieval-dependent QA with SPARQL-verifiable answers, recent work advances reliable long-horizon reasoning along orthogonal axes. Multi-agent methods (zhou2025reso; zhu2025swarm) improve internal coordination and exploration but do not test temporal grounding on evolving facts. RL post-training reveals scaling laws for mathematical reasoning, yet its gains must transfer under recency-sensitive evaluation (tan2025scaling). In multimodal and embodied settings, Visual-O1 and BMMR target ambiguity and cross-discipline reasoning, while BEAR and Code-as-Monitor emphasize verifiable skills and failure detection (ni2024visualo1understandingambiguousinstructions; xi2025bmmrlargescalebilingualmultimodal; qi2025bear; zhou2024code; kang2025viki). RoboRefer strengthens spatial referring via structured reasoning (zhou2025roborefer). VeriGUI contributes stepwise verification for GUI agents (liu2025verigui). LiveSearchBench uniquely couples compositional constraints with fresh, machine-checkable knowledge to assess reasoning beyond parametric memory.